solve the following:
sqrt 3 csc x - 2 cot x = 0 when 0 degrees less then or equal too x less then 360 degrees. Indicate the restriction of the solution.

Answer by Alan3354(69443)

solve the following:
sqrt 3 csc x - 2 cot x = 0 when 0 degrees less then [sic] or equal too [sic] x less then 360 degrees. Indicate the restriction of the solution.
=================
It's less than, not less then.
----
Write it like this:
sqrt(3)csc x - 2 cot x = 0, 0 <= x <= 360
---
sqrt(3)/sin - 2(cos/sin) = 0
Multiply by sine
sqrt(3) - 2cos = 0
cos(x) = sqrt(3)/2
x = 30 degs, 330 degs